# In-Context Learning Strategies Emerge Rationally

### v1

@@ B.2 Broader Work on Understanding ICL and Meta-Learning
A crucial benefit of our Bayesian modeling framework is that one can retrieve posterior probabilities over the predictors the model predictions are being decomposed into. In other words, we can represent the model predictions as an interpolation of the predictors’ outputs (since posterior probabilities sum to 1). Park et al. [26] propose a similar approach: “Linear Interpolation of Algorithms” (LIA). Specifically, LIA represents model predictions as an interpolation of the predictors’ outputs, directly fitting the weights for interpolation. That is, for each condition of training time ($N$), task diversity ($D$) , LIA requires fitting a set of parameters, summing up to a total of $N\times D$ parameters fit (across many of the maps in our work, this would mean over $900$ parameters per context-size and task dimensionality condition). Meanwhile, our framework, by defining a precise functional form that explicates the role of $N$ and $D$, minimizes the number of terms needed to perform fitting to 3—these terms are primarily related to model family, its learning dynamics, and intrinsic randomness of the data, hence making them hard to explicate. Crucially, one can see our framework as providing grounding to LIA: under the hood, LIA is trying to identify the posterior probabilities for all experimental settings! Moreover, our framework offers an explanation for why change in predictor weights occurs during training, and importantly, can predict dynamics of $N,D$ conditions it was not trained on, which LIA cannot do since it has to be fit separately to each $N,D$ condition.
@@
Beyond the papers discussed above, there have been several complementary efforts to understand ICL from different perspectives (see the recent summary by Lampinen et al. [37] for a longer review). For example, several papers analogize ICL as implicitly performing optimization to learn novel tasks [22, 52, 36]; meta-learning in general [79, 80]; develop scaling laws for the sample efficiency of ICL [72]; identify limits of tasks that can be learned in-context [14, 17, 81]; demonstrate sudden learning curves exist in ICL [53, 10]; and characterize mechanisms employed by a model to perform ICL [13, 34, 82, 31]. Broadly, all these papers offer useful insights that are complementary to ours.
@@ Appendix C Future Work
  • Does our Framework Explain other Pretraining? we believe the competition dynamic characterized in our work is similar to the one demonstrated by Qin et al. [85] in a language modeling task. There, the authors show that depending on the diversity of rules of the language and training time, a model can either learn a bag of heuristics or the underlying grammar to generate sentences from the language. It may be possible to explain the phenomenology elicited in that work via the hierarchical Bayes lens we take in this paper, since there likely exists a tradeoff between compressibility and loss of a heuristic vs. grammar-learning solution.
@@ Appendix D Derivations
  • A2: Neural networks exhibit a simplicity bias, such that the complexity of implementing a predictor $Q$ follows a prior of the form $p(Q)=2^{-K(Q)^{\beta}}$, where $K(Q)$ denotes the estimated Kolmogorov complexity of $Q$. To accommodate the fact that we care about implementation in a Transformer, we instantiate a free parameter ($\beta$) on complexity while preserving the form of a universal prior.
@@ D.1 Log-Posterior Odds
$\propto\log\int_{\theta\in\Theta_{Q}}P(S_{\mathcal{T}_{\text{train}}}^{(N)}|\theta)\,P(\theta|S_{\mathcal{T}_{\text{train}}}^{(1)},\dots,S_{\mathcal{T}_{\text{train}}}^{(N-1)},T,H)$
@@
=$\stackrel{{\scriptstyle\text{{A2}}}}{{=}}N_{\text{eff}}\underbrace{\frac{1}{N_{\text{eff}}}\sum_{n=1}^{N_{\text{eff}}}\log P(S_{\mathcal{T}_{\text{train}}}^{(n)}|Q)}_{\text{Log likelihood of data under predictor $Q$}}-K(Q)^{\beta}\cdot\log_{{\rm e}}2$
$=N_{\text{eff}}\,L(Q)-K(Q)^{\beta}\cdot\log_{{\rm e}}2.$ (6)
@@
$=N_{\text{eff}}\,\left(\log\;P(S_{\mathcal{T}_{\text{train}}}^{(n)}|M)-\log\;P(S_{\mathcal{T}_{\text{train}}}^{(n)}|G))\right)+\left(-K(M)^{\beta}+K(G)^{\beta}\right)\cdot\log_{{\rm e}}2$
$=N_{\text{eff}}\,\left(L(G)-L(M)\right)-\left(K(M)^{\beta}-K(G)^{\beta}\right)\cdot\log_{{\rm e}}2$
@@
$=\frac{\sum_{n=1}^{N}(L(n)-L(\infty))\delta n}{L(Q)}$
$=\frac{1}{L(Q)}\sum_{n=1}^{N}\frac{A}{n^{\alpha}}\delta n$
$=\frac{A}{L(Q)}N^{1-\alpha}\int_{0}^{1}\frac{1}{\hat{n}^{\alpha}}\delta\hat{n}$
@@
where $\hat{n}=\nicefrac{{n}}{{N}}$ and $\gamma=\nicefrac{{A}}{{L(Q)}}$ is a constant that subsumes the loss of the predictor $Q$ and the constant $A$ from our assumed form of power-law scaling, which depends on the random loss of a network and effects of hyperparameters like batch-size and sequence lengths used to define the train data.
@@ D.2 Converting from Posterior-Odds to a Predictive Model
$\boxed{h_{\text{pred}}(\bm{s})=\sigma(\eta(N,D))\,M(s)+(1-\sigma(\eta(N,D)))\,G(s).}$ (10)
@@ Appendix E Two-Hypotheses Threshold: Minimum amount of training to enable the Hierarchical Bayesian Model
One can reasonably expect our proposed Hierarchical Bayesian model to explain learning dynamics of in-context learning will not be predictive of a Transformer’s behavior early-on in training, for otherwise we are saying even an untrained model perfectly generalizes. In actuality, the Transformer becomes amenable to approximation by our model after some minimal amount of training has occurred. To automatically calculate whether we have finished this regime of training, we calculate an “optimal” interpolation between the two predictors if they were capable of explaining the model behavior: specifically, we rely on the relative distance as an estimation of the optimal interpolation weighting, and use it as a baseline to compare model outputs with. In particular, we compute the loss between this optimal interpolation baseline and our trained Transformer model, and if this loss is below a threshold, we claim our theoretical model is applicable. We call this threshold the two-hypotheses threshold (see Fig. 6).
To define the two-hypotheses threshold, we make the observation that while the loss between Transformer and interpolating predictor can be large to begin with, it very quickly reduces to a small value. We can expect this latter regime is where our theoretical model is most likely to accurate at modeling the trained Transformer’s behavior. Motivated by this, we heuristically choose the two-hypotheses threshold as a loss value 20% higher than minimum for Balls & Urns and Classification, and 10% higher than minimum for Linear Regression, on the scale defined from minimum to maximum loss (we find that a stricter threshold is required for linear regression to surpass the early high-loss regime, given the larger variance in interpolation loss values. See Fig. 9).
@@ F.1 Training and Model Details
    • We conduct a separate experiment in which we attempt to elicit transience in higher task diversity settings ($D\in\{2^{8},2^{9}\}$). To do so on a reasonable compute budget ($2$M, $10$M steps, respectively), we, similar to prior work [24, 29], have to intervene on the training pipeline. However, unlike prior work that often relies on weight decay for this purpose, we use learning rate annealing and find it to be sufficient. Specifically, we rely on an inverse square root schedule for decaying the learning rate with number of dimensions, context length, and MLP expansion fixed to $8$, $128$, and $4$ respectively. We train up to $D=2^{7}$ for $100$K steps, and then train with $D=2^{8}$ for $2$M steps, and with $D=2^{9}$ steps for $10$M steps. Results of this experiment are shown in Fig. 6(c).
  • Linear Regression. Models of hidden dimension size $64$ are trained for $100$K steps, with $5$K warmup steps, at a constant learning rate of $5\times 10^{-4}$ and batch-size of $128$. For our analysis, we derive experimental settings from combinations of task-dimensionality, which varies in the set $\{8,12,16\}$; context length, which varies in the set $\{16,32,64\}$; and MLP expansion factor, which varies in the set $\{0.5,4,8\}$.
@@ F.2 Analysis Details
For OOD evaluation of both the Transformer and our procedurally defined predictors, i.e., the memorizing predictor $M$ and generalizing predictor $G$, we draw $500$ sequences from $500$ unseen tasks (however, following still the same task distribution $\mathcal{T}_{\text{true}}$). In comparison, ID evaluation involves $500$ sequences from seen tasks. If task-diversity $D$ is less than $500$, sequences from the same task may be seen multiple times. We note that the ID evaluation for classification is slightly different, following the popularly used pipeline of ‘in-weights learning’ first introduced by Chan et al. [20]: herein, one specifically ensures that a copy of the test item does not appear in the context, thus disallowing the use of an in-context copying solution and testing memorization more explicitly.
@@
Recall that relative distance, for a given distance measure $d(.,.)$ between two functions or distributions (e.g., KL-divergence or Euclidean distance), is defined as $d_{\text{rel}}=\nicefrac{{(r+1)}}{{2}}$, where $r:=\frac{d(h,G)-d(h,M)}{d(G,M)}$ and $h(.)$ denotes the Transformer model trained from scratch. This metric implicitly makes the assumption that in some function space, the model $h(.)$ lies on a line between the predictors $M$ and $G$. Correspondingly, for the scenarios this assumption is violated, the value of $d_{\text{rel}}$ can go outside the range 001111. This relatively rarely, but nevertheless noticeably, occurs (e.g., if the model implements the optimal constant solution early on in training for linear regression). Accordingly, we clamp the metric between 001111.
@@
One can reasonably expect our proposed Hierarchical Bayesian model to explain learning dynamics of in-context learning will not be predictive of a Transformer’s behavior early-on in training, for otherwise we are saying even an untrained model perfectly generalizes. In actuality, the Transformer becomes amenable to approximation by our model after some minimal amount of training has occurred. To automatically calculate whether we have finished this regime of training, we calculate an “optimal” interpolation between the two predictors if they were capable of explaining the model behavior: specifically, if the euclidean distance had to be minimum between the interpolation of predictors and Transformer outputs, we can identify a closed-form solution for the interpolation weights and use it as a baseline to compare model outputs with. In particular, we compute the loss between this optimal interpolation baseline and our trained Transformer model, and if this loss is below a threshold, we claim our theoretical model is applicable now. This threshold is defined by simply computing the maximum and minimum loss with respect to the interpolating predictor, and taking the value 20% higher than minimum on the scale defined from minimum to maximum loss.
@@
  • Computation of average log likelihood per predictor. For every experimental condition we first record the token-level log-likelihood that each predictor assigns to the in-distribution sequences, following the procedure described above. Because the irreducible error term cancels when models are compared, we restrict KL-based evaluations to the Balls & Urns and classification tasks, treating the linear-regression setting separately. To summarize performance we need the mean log-likelihood per token, yet the empirical loss distribution is, at times, quite skewed: tokens typically incur near-zero loss, whereas a small fraction of early tokens produce large spikes. A naive arithmetic mean therefore converges slowly and exhibits high variance. Thus, we instead use median of means, an estimator for the true mean that has better convergence under long tailed distributions.
@@ Appendix H Main Results Across All Settings
H.1 Task Diversity Threshold
We find the phenomenon of a task diversity threshold [25] to be very robust across settings and experimental conditions. More specifically, we consistently find that increasing task diversity yields a transition in Transformer behavior from behaving like a memorizing predictor to behaving like a generalizing predictor. In the following, we present evidence of this phenomenon for ID sequences. See results in following pages.
@@ H.1.1 Balls & Urns
Figure 14: Task Diversity Threshold across Balls & Urns conditions. Red dashed line indicates the memorizing solution $M$, blue dashed line indicates the generalizing solution $G$, and black solid line indicates Transformer behavior at the end of training ($100$K steps).

### v2

@@ B.3 Broader Work on Understanding ICL
A crucial benefit of our Bayesian modeling framework is that one can retrieve posterior probabilities over the predictors the model predictions are being decomposed into. In other words, we can represent the model predictions as an interpolation of the predictors’ outputs (since posterior probabilities sum to 1). Park et al. [26] propose a similar approach: “Linear Interpolation of Algorithms” (LIA). Specifically, LIA represents model predictions as an interpolation of the predictors’ outputs, directly fitting the weights for interpolation. That is, for each condition of training time ($N$), task diversity ($D$), LIA requires fitting a set of parameters, summing up to a total of $N\times D$ parameters fit (across many of the maps in our work, this would mean over $900$ parameters per combination of context-length, task dimensionality, and MLP width). Meanwhile, our framework, by defining a precise functional form that explicates the role of $N$ and $D$, minimizes the number of terms needed to perform fitting to 3—these terms are primarily related to model family, its learning dynamics, and intrinsic randomness of the data, hence making them hard to explicate. Crucially, one can see our framework as providing grounding to LIA: under the hood, LIA is trying to identify the posterior probabilities for all experimental settings! Moreover, our framework offers an explanation for why change in predictor weights occurs during training, and importantly, can predict dynamics of $N,D$ conditions it was not trained on, which LIA cannot do since it has to be fit separately to each $N,D$ condition.
@@
Beyond the papers discussed above, there have been several complementary efforts to understand ICL from different perspectives (see the recent summary by Lampinen et al. [37] for a longer review). For example, several works, some briefly discussed above, characterized the trade-off between memorizing and generalizing in a more mechanistic way, as a competition dynamic between circuits [29, 33, 24, 26]. We see our results as elucidating the forces driving this competition, as well as demonstrating its consistency across several settings. Additionally, several papers analogize ICL as implicitly performing optimization to learn novel tasks [22, 53, 36] or to meta-learning in general [91]; develop scaling laws for the sample efficiency of ICL [73]; identify limits of tasks that can be learned in-context [14, 17, 92]; demonstrate sudden learning curves exist in ICL [54, 10]; and characterize mechanisms employed by a model to perform ICL [13, 34, 93, 31]. Broadly, all these papers offer useful insights that are complementary to ours.
@@ Appendix C Future Work
  • Does our Framework Explain other Pretraining Phenomena? we believe the competition dynamic characterized in our work is similar to the one demonstrated by Qin et al. [96] in a language modeling task. There, the authors show that depending on the data diversity and complexity, a model can either learn a bag of heuristics or the underlying grammar to generate sentences from the language. It may be possible to explain the phenomenology elicited in that work via the hierarchical Bayes lens we take in this paper, since there likely exists a tradeoff between compressibility and loss of a heuristic vs. grammar-learning solution.
@@ Appendix D Derivations
  • A2: Neural networks exhibit a bias toward simpler solutions. Specifically, using $K(Q)$ to denote the Kolmogorov complexity for predictor $Q$, we accommodate the Transformer-specific implementation cost by defining $\tilde{K}_{\text{T}}(Q)=K(Q)^{\beta}$. Then, taking the form of a universal prior, the prior probability of learning a predictor $Q$ is $p(Q)\;\propto\;2^{-\tilde{K}_{\text{T}}(Q)}=2^{-K(Q)^{\beta}}$.
@@ D.1 Log-Posterior Odds
$\propto\log\int_{\theta\in\Theta_{Q}}P(S_{\mathcal{T}_{\text{train}}}^{(1)},\dots,S_{\mathcal{T}_{\text{train}}}^{(N)}|\theta)\,P(\theta|T,H)$
@@
$\stackrel{{\scriptstyle\text{{A2}}}}{{\approx}}N_{\text{eff}}\underbrace{\frac{1}{N_{\text{eff}}}\sum_{n=1}^{N_{\text{eff}}}\log P(S_{\mathcal{T}_{\text{train}}}^{(n)}|Q)}_{\text{Average log likelihood of data under predictor $Q$}}-K(Q)^{\beta}\cdot\log_{{\rm e}}2$
$=-N_{\text{eff}}\,\overline{L_{Q}}(S_{\mathcal{T}_{\text{train}}})-K(Q)^{\beta}\cdot\log_{{\rm e}}2.$ (6)
@@
$=N_{\text{eff}}\,\left(\overline{L_{G}}(S_{\mathcal{T}_{\text{train}}}(D))-\overline{L_{M}}(S_{\mathcal{T}_{\text{train}}}(D))\right)-\left(K(M)^{\beta}-K(G)^{\beta}\right)\cdot\log_{{\rm e}}2$
@@
$=\frac{\sum_{n=1}^{N}(L(n)-L(\infty))\delta n}{\overline{L_{Q}}}$
$=\frac{1}{\overline{L_{Q}}}\sum_{n=1}^{N}\frac{A}{n^{\alpha}}\delta n$
$=\frac{A}{\overline{L_{Q}}}N^{1-\alpha}\int_{0}^{1}\frac{1}{\hat{n}^{\alpha}}\delta\hat{n}$
@@
where $\overline{L_{Q}}=\overline{L_{Q}}(S_{\mathcal{T}_{\text{train}}}(D))$, $\hat{n}=\nicefrac{{n}}{{N}}$ and $\gamma=\nicefrac{{A}}{{\overline{L_{Q}}}}$ is a constant that subsumes the loss of the predictor $Q$ and the constant $A$ from our assumed form of power-law scaling, which depends on the random loss of a network and effects of hyperparameters like batch-size and sequence lengths used to define the train data.
@@ D.2 Converting from Posterior-Odds to a Predictive Model
$\boxed{h_{\text{pred}}(\bm{s})=\sigma(\eta(N,D))\,M(\bm{s})+(1-\sigma(\eta(N,D)))\,G(\bm{s}).}$ (10)
@@ Appendix E Two-Hypotheses Threshold: Minimum amount of training to enable the Hierarchical Bayesian Model
One can reasonably expect our proposed Hierarchical Bayesian model to explain learning dynamics of in-context learning will not be predictive of a Transformer’s behavior early-on in training, for otherwise we are saying even an untrained model perfectly generalizes. In actuality, the Transformer becomes amenable to approximation by our model after some minimal amount of training has occurred. To automatically calculate whether we have finished this regime of training, we calculate an “optimal” interpolation between the two predictors if they were capable of explaining the model behavior: specifically, we rely on the relative distance as an estimate of the optimal interpolation weighting, and use it as a baseline to compare model outputs with. In particular, we compute the loss between this optimal interpolation baseline and our trained Transformer model, and if this loss is below a threshold, we claim our theoretical model is applicable. We call this threshold the two-hypotheses threshold (see Fig. 6).
To define the two-hypotheses threshold, we make the observation that while the loss between Transformer and interpolating predictor can be large to begin with, it very quickly reduces to a small value. We can expect this latter regime is where our theoretical model is most likely to accurate at modeling the trained Transformer’s behavior. Motivated by this, we heuristically choose the two-hypotheses threshold as a loss value 20% higher than minimum for Balls & Urns and Classification, and 10% higher than minimum for Linear Regression, on the scale defined from minimum to maximum loss (we find that a stricter threshold is required for linear regression to surpass the early high-loss regime, given the larger variance in interpolation loss values).
@@ F.1 Training and Model Details
    • We conduct a separate experiment in which we attempt to elicit transience in higher task diversity settings ($D\in\{2^{8},2^{9}\}$). To do so on a reasonable compute budget ($2$M, $10$M steps, respectively), we, similar to prior work [24, 29], have to intervene on the training pipeline. However, unlike prior work that often relies on weight decay for this purpose, we use learning rate annealing and find it to be sufficient. Specifically, we rely on an inverse square root schedule for decaying the learning rate with number of dimensions, context length, and MLP expansion fixed to $8$, $128$, and $4$ respectively. We train up to $D=2^{7}$ for $100$K steps, and then train with $D=2^{8}$ for $2$M steps, and with $D=2^{9}$ steps for $10$M steps. Results of this experiment are shown in Fig. 6(c) and App. K.
  • Linear Regression. Models of hidden dimension size $64$ are trained for $100$K steps, with $5$K warmup steps, at a constant learning rate of $5\times 10^{-4}$ and batch-size of $128$. For our analysis, we derive experimental settings from combinations of task-dimensionality, which varies in the set $\{8,12,16\}$; context length, which varies in the set $\{16,32,64\}$; and MLP expansion factor, which varies in the set $\{0.5,4,8\}$. Like in Balls & Urns, we additionally train a setting using learning rate annealing (setting parameters are task dimensionality of $8$, Context length of $16$, and MLP expansion factor of $4$. We train once with warmup of $500$, and once with warmup of $5000$. See results in app. K).
@@ F.2 Analysis Details
For OOD evaluation of both the Transformer and our procedurally defined predictors, i.e., the memorizing predictor $M$ and generalizing predictor $G$, we draw $500$ sequences from $500$ unseen tasks (however, following still the same task distribution $\mathcal{T}_{\text{true}}$). In comparison, ID evaluation involves $500$ sequences from seen tasks. If task-diversity $D$ is less than $500$, sequences from the same task may be seen multiple times. We note that in addition to the standard ID evaluation, there is another method for ID evaluation in Classification that is slightly different. Following the popularly used pipeline of ‘in-weights learning’ (IWL) first introduced by Chan et al. [20], one specifically ensures that a copy of the test item does not appear in the context, thus making the use of an in-context copying solution ineffective, and testing memorization more explicitly.
Fig. 1 details.
In this figure, we use context length of 128, MLP expansion factor of 4, and task dimensionality of 8 for Balls & Urns. Linear Regression uses similar parameters, only with a context length of 32, and Classification uses similar parameters, only with MLP expansion factor of 8. In all plots displaying task diversity effects, we hold $N=100$K. For classification, we show the IWL method for ID evaluation. In the transience figures, we use $D=256$ for Balls & Urns, $D=64$ for Linear Regression, and $D=512$ for Classification. For the comparison of relative distance maps, we show maps from the Balls & Urns setting with the parameters described above.
@@
Recall that relative distance, for a given distance measure $d(.,.)$ between two functions or distributions (symmetrized KL-divergence or Euclidean distance), is defined as $d_{\text{rel}}=\nicefrac{{(r+1)}}{{2}}$, where $r:=\frac{d(h,G)-d(h,M)}{d(G,M)}$ and $h(.)$ denotes the Transformer model trained from scratch. This metric implicitly makes the assumption that in some function space, the model $h(.)$ lies on a line between the predictors $M$ and $G$. Correspondingly, for the scenarios this assumption is violated, the value of $d_{\text{rel}}$ can go outside the range 001111. This occurs relatively rarely, but nevertheless noticeably (e.g., if the model implements the optimal constant solution early on in training for linear regression). Accordingly, we clamp the metric between 001111.
@@
One can reasonably expect our proposed Hierarchical Bayesian model to explain learning dynamics of in-context learning will not be predictive of a Transformer’s behavior early on in training, for otherwise, we are saying even an untrained model perfectly generalizes. In actuality, the Transformer becomes amenable to approximation by our model after some minimal amount of training has occurred. To automatically calculate whether we have finished this regime of training, we use the relative distance as an estimate of an “optimal” interpolation weight between the two predictors. We compute the loss between this optimal interpolation baseline and our trained Transformer model, and if this loss is below a threshold, we claim our theoretical model is applicable now (see details on our choice of threshold in App. E)
@@
  • Computation of average log likelihood per predictor. For every experimental condition we first compute the token-level log-likelihood that each predictor assigns to the in-distribution sequences (note that we use standard ID sequences for classification, not the IWL sequences). Because the irreducible error term cancels when models are compared, we use KL-based evaluations to the Balls & Urns and classification tasks, treating the linear-regression setting separately. To summarize performance, we need the mean log-likelihood per token, yet the empirical loss distribution is, at times, quite skewed: most tokens later in the context incur near-zero loss, whereas a small fraction of early tokens produce large spikes. Therefore, a naive arithmetic mean converges slowly and exhibits high variance. Thus, we instead use median of means, an estimator for the true mean that has better convergence under long-tailed distributions.
@@
Novel Predictions Analysis Details.
To show sub-linear sample efficiency and a sigmoidal curve in $N^{1-\alpha}$, we fit a parameterized logistic $\frac{a}{1+\exp(-b(N^{1-\alpha}-N_{0}))}$ with free parameters $a,b,N_{0}$ to each training run (constant $D$ value), via scipy curvefit function. We use the $\alpha$ value given by the Bayesian model. Curve fits are shown in Fig. 6(a). To compute the second derivative of the relative distance (Fig. 6), we simply use parameters for the logistic fits described above (which provide very close fits, as can be seen in Fig. 6(a)), then compute the second derivative based on the form for the second derivative of a parametrized logistic. To plot the vector field, we normalize both $U$ and $V$ directions by the larger value among the 90th percentile values for $U$ and $V$.
@@ Appendix H Main Results Across All Settings
H.1 Task Diversity Effects
We find task diversity effects [25] to be very robust across settings and experimental conditions. More specifically, we consistently find that increasing task diversity yields a transition in Transformer behavior from behaving like a memorizing predictor to behaving like a generalizing predictor. In the following, we present evidence of this phenomenon for ID sequences. See results in following pages.
@@ H.1.1 Balls & Urns
Figure 14: Task Diversity Effects Across Balls & Urns Conditions. Red dashed line indicates the memorizing solution $M$, blue dashed line indicates the generalizing solution $G$, and black solid line indicates Transformer behavior at the end of training ($100$K steps).